OpenStudy (anonymous):

how many solutions are there for 5x^2+7x-4=0? a) 0 b) 1 c) 2 d) 3

OpenStudy (anonymous):

have you tried to solve it ? :)

OpenStudy (anonymous):

well i'm in the process. ;)

OpenStudy (anonymous):

what did you get? :) you have 3 cases \[\LARGE b^2-4ac>0\]-Two solutions \[\LARGE b^2-4ac=0\]-One solution \[\LARGE b^2-4ac<0\]-No solution... (Eliminate option "d" now) .. ;)

OpenStudy (anonymous):

\[\LARGE \sqrt{b^2-4ac}=\sqrt{7^2-4\cdot 5\cdot (-4)}=...?\]

OpenStudy (anonymous):

\[\LARGE \sqrt{49-20\cdot (-4)}=\sqrt{49+80}=... \] because \[\LARGE (-)\cdot (-)=(+)\] and 49+80=129 so \[\LARGE \Delta>0\]----- Two solutions... answer is "C"

OpenStudy (anonymous):

omg i was halfway right. ugh, im so proud. :') thanks. lol, you should be my tutor.

OpenStudy (anonymous):

lol, Glad to help. I'm sure you already have a tutor (not me) , your professor . Good Luck!

OpenStudy (anonymous):

lol, she sucks. but thanks anyways. :p